# Continuous-Time Neural Networks Can Stably Memorize Random Spike Trains

### v1

@@ Abstract
Index Terms:
Spiking neural networks, stable memorization
I Introduction
Biological neural networks work in continuous time, without a clock, and with (apparently) imprecise and noisy neurons [1]. It is thus not obvious how memories can be stably represented at all. Nonetheless, biological neural networks appear to have impressive storage capacity and the ability to control muscles with high temporal precision. Understanding this better is of interest in its own right, and it may give ideas for designing future neuromorphic hardware [2, 3, 4, 5].
It is an old idea that memories can be encoded as stable attractors of the network dynamics [6, 7, 8]. The classical Hopfield networks [9] works with point attractors, and cyclic attractors have been proposed for representing repetitive biological behaviors such as breathing, running, or flying [10]. A related concept from the literature are polychronous groups of neurons as observed by Izhikevich [11].
We also note that many studies of neural dynamics (of continuous-time recurrent networks) have focused on networks in which the weights are randomly assigned (i.e., not learned or precomputed), and this setting has been observed to promote chaotic behavior [12].
@@ I Introduction
Figure 1: A score of spike trains $x_{1},...,x_{L}$ produced by $L$ neurons, with minimum delay $\tau_{0}$ (= distance between the tickmarks) between spikes.
@@
We will work with (a variation of) a standard continuous-time neural network model with random transmission delays. The latter are essential; we thus reconfirm the observation from [11] that axonal delays are not a nuisance, but a valuable resource.
@@
The paper is structured as follows. The network model is defined in Section II. The precise problem statement and the main experimental results are given in Section III. The computation of the synaptic weights and further experimental results are described in Section IV. The capacity of such networks is discussed in Section V. In Section VI, the results of Section III are complemented with an eigenvalue analysis of a linearized version of the network model. Associative recall is demonstrated in Section VII, and the conclusions are offered in Section VIII. Some technical details are given in the Appendices. The software for the experiments in this paper is available on GitHub [^1] https://github.com/haguettaz/RSNN.
II Spiking Neural Network Model
II-A Spike Trains and Network Structure
@@
II-B The Spiking-Neuron Model
We will use a version of the Spike Response Model (SRM) [13, 14], which offers just enough biological (or physical) realism to be useful for our purpose.
@@ II-B The Spiking-Neuron Model
The pivotal quantity of each neuron is its potential $z_{\ell}(t)$: the neuron fires (i.e., it produces a spike) whenever $z_{\ell}(t)$ crosses a threshold $\theta_{\ell}(t)$, as will be detailed in Sec. II-C.
@@
II-C Refractory Period and Threshold Noise
@@
All neurons have the same fixed nominal threshold $\theta_{0}>0$. However, neuronal computation is subject to noise at various levels [15], which manifests in shifted, false, or missed firings. In this paper, the primary way to model such noise is to work with a random threshold $\theta_{\ell}(t)$ as follows: after each firing of neuron $\ell$, its threshold $\theta_{\ell}$ (for the next firing) is newly (and independently) sampled from a Gaussian distribution with mean $\theta_{0}$ and variance $\sigma_{\theta}^{2}$, as illustrated in Fig. 4.
@@
III Problem Statement and Main Result
@@
III-A Just Standard Supervised Learning?
Obviously, the stated problem may be viewed as a form of supervised learning, where each neuron is trained to predict its next prescribed firing. However, standard methods of supervised learning [16] will not enable the network to stably reproduce the prescribed score of spike trains autonomously: small errors in the firing times will quickly accumulate and the temporal coherence of the spike score will be lost (cf. the first row of Table III and Table VI).
@@ III-A Just Standard Supervised Learning?
III-B The Prescribed Score of Spikes: Periodic and Random
In order to simplify the discussion, the prescribed score of spike trains is periodic with period $T\gg\tau_{0}$. In other words, we wish the memorized spike score to be a cyclic attractor. (The generalization to multiple cyclic attractors is obvious and will be demonstrated in Sec. VII.)
We will consider random periodic spike scores, which we wish the network to reproduce with probability close to one. In fact, it is easy to see that there exist spike scores that the network cannot be made to reproduce [^2] E.g., a neuron cannot fire if it receives no firings at any of its synapses., but we wish the probability of all such unfeasible spike scores to be close to zero.
The $L$ random spike trains (one for each neuron) are generated independently. The probability law of each spike train is a suitably modified Poisson process with spikes separated by at least $\tau_{0}$. Both the exact definition and the practical generation (i.e., the sampling) of such random spike trains are detailed in Appendix B. In particular, the expected number of spikes (of each neuron, per period $T$) as a function of the firing rate $\lambda$ is given by (B.13) and plotted in Fig. 11.
III-C Measuring Accuracy and Stability
@@
Clearly, (4) and (5) coincide if $|\mathcal{S}_{\ell}^{t_{0}}|=|\breve{\mathcal{S}}_{\ell}^{0}|$; in this case, low values of $\operatorname{pr}(t_{0})$ and $\operatorname{rc}(t_{0})$ indicate strong firing jitter. The condition $\operatorname{pr}(t_{0})\gg\operatorname{rc}(t_{0})$ indicates many missing firings while $\operatorname{pr}(t_{0})\ll\operatorname{rc}(t_{0})$ indicates many false firings, cf. Sec.III-E and Sec.VII.
@@
III-D Drift and Very-Long-Term Stability
Any two physical clocks that are not synchronized will inevitably drift away from each other (unless there is absolutely no noise). This applies also to the problem statement in this section: the threshold noise may have a net effect of pushing all spikes forward (or backward) in time, which is unrecoverable (see also Sec. VI). Therefore, our measures of precision (4) and recall (5) cover only a single period rather than the whole time axis.
Moreover, the reader may have noticed that, with threshold noise as in Sec. II-C, the experimentally observed temporal stability (cf. Sec. III-E) cannot last forever. For example, it is possible that the momentary thresholds $\theta_{\ell}(t)$ (i.e., $\theta_{0}$ plus threshold noise) of all neurons are set to such high values that all firings are suppressed, which stalls the network in that state. However, in some range of parameters, the probability of such derailments is so small that they are never observed in numerical experiments.
III-E Experimental Results
TABLE I: Accuracy of autonomous reproduction with threshold noise. Every row summarizes 100 repetitions of the same experiment.
@@ III-E Experimental Results
TABLE II: Default values of the model parameters
in all numerical experiments.
@@
$L$ number of neurons (Sec. II-A) $200$
$K$ number of inputs per neuron (Sec. II-A) $500$
$\beta$ input kernel parameter (Sec. II-B) $\tau_{0}$
$d_{\text{min}}$ minimal axonal delay (Sec. II-B) $0.1\,\tau_{0}$
$d_{\text{max}}$ maximal axonal delay (Sec. II-B) $10\,\tau_{0}$
$\theta_{0}$ nominal threshold (Sec. II-C) $1$
$\sigma_{\theta}$ standard deviation of threshold noise (Sec. II-C) $0.1\,\theta_{0}$
$T$ period of firing score (Sec. III-B and App. B) $50\,\tau_{0}$
$\lambda$ random-firing rate (App. B) $0.2/\tau_{0}$
@@
Some exemplary experimental results are reported in Table I, which shows the precision (4) and the recall (5) for different values of $L$ (the number of neurons) and $\sigma_{\theta}$ (the standard deviation of the threshold noise). The values of all other parameters are listed in Table II, which shows the default values of all parameters in all numerical experiments of this paper. With $\lambda$ and $T$ as in Table II, the expected total number of spikes (of all neurons, during one period of duration $T$) is about $7L$ (cf. Fig. 11 in Appendix B).
In these experiments, the synaptic weights are computed as described in Sec. IV (below). The network is initialized with the exact correct firing times; then it is left to operate autonomously (with threshold noise as in Section II-C). After about five periods, the precision and the recall have reached a steady state, which is reported in Table I (after 20 periods); running the simulations longer does not change anything (but see Sec. III-D for extreme time scales).
Every row of Table I summarizes 100 repetitions of the same experiment, each with a new random network and a new random prescribed firing score; every row reports the median, the minimum, and the maximum values of (4) and (5) of all 100 repetitions.
@@
IV Computing the Synaptic Weights
@@
$z_{\ell}(t)<\theta_{\mathrm{r}}$ unless $s-\varepsilon_{\mathrm{s}}<t<s+\tau_{0}$ (10)
for some $s\in\breve{S}_{\ell}$
@@
Similar conditions for the potential $z_{\ell}(t)$ have been used in [17], which, however, did not address temporal stability. The minimal-slope condition (11) appears to be a new ingredient, which is essential for temporal stability; the effect of this condition is demonstrated in Table III (see also Table VI).
@@
with $\nu=1$ (resulting in sparse weights) or $\nu=2$, subject to (8)–(12). The effect of this additional regularization is demonstrated in Table IV (see also Table VI).
@@
TABLE III: Effect of the minimum-slope condition (11). Every row summarizes 100 repetitions of the same experiment.
@@
TABLE IV: Effect of the additional regularization by (13). Every row summarizes 100 repetitions of the same experiment.
@@
From the experiments in Tables III and IV, and from many other experiments, we observe that satisfying (8)–(12) (with sufficiently large $\dot{\theta}_{\mathrm{s}}$ and sufficiently small $w_{\mathrm{b}}$ and $\theta_{\mathrm{r}}$, for all neurons simultaneously) suffices to guarantee temporal stability at some level of threshold noise. The feasibility of (8)–(12) can therefore be used as a proxy for stable memorization.
V Memory Capacity
Clearly, for any fixed network, there must be an upper limit on the maximal period $T$ of the memorizable content. Some pertinent (and typical) experimental results are reported in Fig. 8. In these experiments, the feasibility of (8)–(12) is used as a proxy for stable memorization (as explained above). Fig. 8 shows the probability of this feasibility as a function of $T$ for different values of $K$, with other parameters as in Table II.
@@ V Memory Capacity
Figure 8: Logistic regression (with squared error fitting) of the empirical probability of feasibility vs. the period $T$ (with relevant parameters as in Table II).
@@
Note that Fig. 8 suggests that the maximal length $T$ of stably memorizable content grows at least linearly with $K$.
VI Small-Jitter Analysis
A necessary condition for temporal stability is that the propagation of spike jitter through the network is damped rather than increased. For small jitter, this can be analyzed by the eigenvalues of a linearized model, as was done in [12]. We are now going to do this as well, and we will find that satisfaction of (8)–(12) (with sufficiently large $\dot{\theta}_{\mathrm{s}}$ and sufficiently small $w_{\mathrm{b}}$ and $\theta_{\mathrm{r}}$) guarantees the suppression of sufficiently small jitter, in agreement with the pertinent remarks in Sec. IV.
VI-A The Math
@@
where we assume (without loss of essential generality) that the influence of $\Delta_{n-n^{\prime}}$ for $n^{\prime}>N$ can be neglected. The derivatives (15) are given by (A.9) in Appendix A.
@@
VI-B The Use
The eigenvalue condition (20) is necessary, but not sufficient, for temporal stability (with finite threshold noise variance) in the sense of Sec. III. However, computing the eigenvalues of $\boldsymbol{\Phi}_{N}$ is less time-consuming than, and provides complementary insight to, simulations as in Sections IIIV.
Complementing the observations in Sec. IV, we experimentally find that satisfying (8)–(12) suffices to satisfy also the eigenvalue condition (20). With the same parameters as for Tables III and IV, we evaluate the eigenvalues of $\boldsymbol{\Phi}_{N}$ and report the two largest (in magnitude) $|\varphi_{1}|$ and $|\varphi_{2}|$ in Tables VI and VI. Every row summarizes 100 repetitions of the same experiment, each with a new random network and a new random prescribed firing score; every row reports the median, the minimum, and the maximum (in magnitude) of these eigenvalues.
Note that the first row of Table VI demonstrates that dropping the minimal-slope condition (11) makes the network unstable.
TABLE V: Effect of the minimum-slope condition (11) on the eigenvalue condition (20), cf. Table III. Every row summarizes 100 repetitions of the same experiment.

### v2

@@ Abstract
1 Introduction
Biological neural networks work in continuous time, without a clock, and with (apparently) imprecise and noisy neurons (Kandel et al., 2021). It is thus not obvious how memories can be stably represented at all. Nonetheless, biological neural networks appear to have impressive storage capacity and the ability to control muscles with high temporal precision. Understanding this better is of interest in its own right, and it may give ideas for designing future neuromorphic hardware (Indiveri & Liu, 2015; Grimaldi et al., 2023; Rubino et al., 2023; Dalgaty et al., 2024).
It is an old idea that memories can be encoded as stable attractors of the network dynamics (Amit, 1989; Poucet & Save, 2005; Ramsauer et al., 2021). The classical Hopfield networks (Hopfield, 1982) works with point attractors, and cyclic attractors have been proposed for representing repetitive biological behaviors such as breathing, running, or flying (Eliasmith, 2005). A related concept from the literature are polychronous groups of neurons as observed by Izhikevich (2006).
We also note that many studies of neural dynamics (of continuous-time recurrent networks) have focused on networks in which the weights are randomly assigned (i.e., not learned or precomputed), and this setting has been observed to promote chaotic behavior (Banerjee et al., 2008).
@@ 1 Introduction
Figure 1: A score of spike trains $x_{1},...,x_{L}$ produced by $L$ neurons, with minimum separation $\tau_{0}$ (= distance between the tickmarks) between spikes.
@@
We will work with (a variation of) a standard continuous-time neural network model with random transmission delays. The latter are essential; we thus reconfirm the observation from Izhikevich (2006) that axonal delays are not a nuisance, but a valuable resource.
@@
The paper is structured as follows. The network model is defined in Section 2. The precise problem statement and the main experimental results are given in Section 3. The computation of the synaptic weights and further experimental results are described in Section 4. The capacity of such networks is discussed in Section 5. In Section 6, the results of Section 3 are complemented with an eigenvalue analysis of a linearized version of the network model. Associative recall is demonstrated in Section 7, and the conclusions are offered in Section Conclusion. Some technical details are given in the Appendices. The software for the experiments in this paper is available on GitHub [^1] https://github.com/haguettaz/RSNN.
2 Spiking Neural Network Model
2.1 Spike Trains and Network Structure
@@
2.2 The Spiking-Neuron Model
We will use a version of the Spike Response Model (SRM) (Gerstner, 1995; Kistler et al., 1997), which offers just enough biological (or physical) realism to be useful for our purpose.
@@ 2.2 The Spiking-Neuron Model
The pivotal quantity of each neuron is its potential $z_{\ell}(t)$: the neuron fires (i.e., it produces a spike) whenever $z_{\ell}(t)$ crosses a threshold $\theta_{\ell}(t)$, as will be detailed in Sec. 2.3.
@@
2.3 Refractory Period and Threshold Noise
@@
All neurons have the same fixed nominal threshold $\theta_{0}>0$. However, neuronal computation is subject to noise at various levels (Faisal et al., 2008), which manifests in shifted, false, or missed firings. In this paper, the primary way to model such noise is to work with a random threshold $\theta_{\ell}(t)$ as follows: after each firing of neuron $\ell$, its threshold $\theta_{\ell}$ (for the next firing) is newly (and independently) sampled from a Gaussian distribution with mean $\theta_{0}$ and variance $\sigma_{\theta}^{2}$, as illustrated in Fig. 4.
@@
3 Problem Statement and Main Result
@@
3.1 Just Standard Supervised Learning?
Obviously, the stated problem may be viewed as a form of supervised learning, where each neuron is trained to predict its next prescribed firing. However, standard methods of supervised learning (Wang et al., 2020) will not enable the network to stably reproduce the prescribed score of spike trains autonomously: small errors in the firing times will quickly accumulate and the temporal coherence of the spike score will be lost (cf. the first row of Table 3 and Table 6).
@@ 3.1 Just Standard Supervised Learning?
3.2 The Prescribed Score of Spikes: Periodic and Random
In order to simplify the discussion, the prescribed score of spike trains is periodic with period $T\gg\tau_{0}$. In other words, we wish the memorized spike score to be a cyclic attractor. (The generalization to multiple cyclic attractors is obvious and will be demonstrated in Sec. 7.)
We will consider random periodic spike scores, which we wish the network to reproduce with probability close to one. In fact, it is easy to see that there exist spike scores that the network cannot be made to reproduce [^2] E.g., a neuron cannot fire if it receives no firings at any of its synapses. Or a neuron sees the same past inputs twice, but is supposed to fire only in one of the two cases., but we wish the probability of all such unfeasible spike scores to be close to zero.
The $L$ random spike trains (one for each neuron) are generated independently. The probability law of each spike train is a suitably modified Poisson process with spikes separated by at least $\tau_{0}$. Both the exact definition and the practical generation (i.e., the sampling) of such random spike trains are detailed in Appendix Appendix B: The Random Periodic Spike Trains. In particular, the expected number of spikes (of each neuron, per period $T$) as a function of the firing rate $\lambda$ is given by (B.13) and plotted in Fig. 11.
3.3 Measuring Accuracy and Stability
@@
Clearly, (4) and (5) coincide if $|\mathcal{S}_{\ell}^{t_{0}}|=|\breve{\mathcal{S}}_{\ell}^{0}|$; in this case, low values of $\operatorname{pr}(t_{0})$ and $\operatorname{rc}(t_{0})$ indicate strong firing jitter. The condition $\operatorname{pr}(t_{0})\gg\operatorname{rc}(t_{0})$ indicates many missing firings while $\operatorname{pr}(t_{0})\ll\operatorname{rc}(t_{0})$ indicates many false firings, cf. Sec.3.5 and Sec.7.
@@
3.4 Drift and Very-Long-Term Stability
Any two physical clocks that are not synchronized will inevitably drift away from each other (unless there is absolutely no noise). This applies also to the problem statement in this section: the threshold noise may have a net effect of pushing all spikes forward (or backward) in time, which is unrecoverable (see also Sec. 6). Therefore, our measures of precision (4) and recall (5) cover only a single period rather than the whole time axis.
Moreover, the reader may have noticed that, with threshold noise as in Sec. 2.3, the experimentally observed temporal stability (cf. Sec. 3.5) cannot last forever. For example, it is possible that the momentary thresholds $\theta_{\ell}(t)$ (i.e., $\theta_{0}$ plus threshold noise) of all neurons are set to such high values that all firings are suppressed, which stalls the network in that state. However, in some range of parameters, the probability of such derailments is so small that they are never observed in numerical experiments.
3.5 Experimental Results
Table 1: Accuracy of autonomous reproduction with threshold noise. Every row summarizes 100 repetitions of the same experiment.
@@ 3.5 Experimental Results
Table 2: Default values of the model parameters in all numerical experiments.
@@
$L$ number of neurons (Sec. 2.1) $200$
$K$ number of inputs per neuron (Sec. 2.1) $500$
$\beta$ input kernel parameter (Sec. 2.2) $\tau_{0}$
$d_{\text{min}}$ minimal axonal delay (Sec. 2.2) $0.1\,\tau_{0}$
$d_{\text{max}}$ maximal axonal delay (Sec. 2.2) $10\,\tau_{0}$
$\theta_{0}$ nominal threshold (Sec. 2.3) $1$
$\sigma_{\theta}$ standard deviation of threshold noise (Sec. 2.3) $0.1\,\theta_{0}$
$T$ period of firing score (Sec. 3.2 and App. Appendix B: The Random Periodic Spike Trains) $50\,\tau_{0}$
$\lambda$ random-firing rate (App. Appendix B: The Random Periodic Spike Trains) $0.2/\tau_{0}$
@@
Some exemplary experimental results are reported in Table 1, which shows the precision (4) and the recall (5) for different values of $L$ (the number of neurons) and $\sigma_{\theta}$ (the standard deviation of the threshold noise). The values of all other parameters are listed in Table 2, which shows the default values of all parameters in all numerical experiments of this paper. With $\lambda$ and $T$ as in Table 2, the expected total number of spikes (of all neurons, during one period of duration $T$) is about $7L$ (cf. Fig. 11 in Appendix Appendix B: The Random Periodic Spike Trains).
In these experiments, the synaptic weights are computed as described in Sec. 4 (below). The network is initialized with the exact correct firing times; then it is left to operate autonomously (with threshold noise as in Section 2.3). After about five periods, the precision and the recall have reached a steady state, which is reported in Table 1 (after 20 periods); running the simulations longer does not change anything (but see Sec. 3.4 for extreme time scales).
Every row of Table 1 summarizes 100 repetitions of the same experiment, each with a new random network and a new random prescribed firing score; every row reports the median, the minimum, and the maximum values of (4) and (5) of all 100 repetitions.
@@
4 Computing the Synaptic Weights
@@
$z_{\ell}(t)<\theta_{\mathrm{r}}$ unless $s-\varepsilon_{\mathrm{s}}<t<s+\tau_{0}$ for some $s\in\breve{S}_{\ell}$ (10)
@@
Similar conditions for the potential $z_{\ell}(t)$ have been used in (Lee et al., 2017), which, however, did not address temporal stability. The minimal-slope condition (11) appears to be a new ingredient, which is essential for temporal stability; the effect of this condition is demonstrated in Table 3 (see also Table 6).
@@
with $\nu=1$ (resulting in sparse weights) or $\nu=2$, subject to (8)–(12). The effect of this additional regularization is demonstrated in Table 4 (see also Table 6).
@@
Table 3: Effect of the minimum-slope condition (11). Every row summarizes 100 repetitions of the same experiment.
@@
Table 4: Effect of the additional regularization by (13). Every row summarizes 100 repetitions of the same experiment.
@@
From the experiments in Tables 3 and 4, and from many other experiments, we observe that satisfying (8)–(12) (with sufficiently large $\dot{\theta}_{\mathrm{s}}$ and sufficiently small $w_{\mathrm{b}}$ and $\theta_{\mathrm{r}}$, for all neurons simultaneously) suffices to guarantee temporal stability at some level of threshold noise. The feasibility of (8)–(12) can therefore be used as a proxy for stable memorization.
5 Memory Capacity
Clearly, for any fixed network, there must be an upper limit on the maximal period $T$ of the memorizable content. Some pertinent (and typical) experimental results are reported in Fig. 8. In these experiments, the feasibility of (8)–(12) is used as a proxy for stable memorization (as explained above). Fig. 8 shows the probability of this feasibility as a function of $T$ for different values of $K$, with other parameters as in Table 2. With Poisson parameter $\lambda=0.2/\tau_{0}$, every neuron fires about 15 times per $100\,\tau_{0}$ (cf. Fig. 11).
@@ 5 Memory Capacity
Figure 8: Empirical probability of feasibility vs. the period $T$, over 100 repetitions (each with a new random network and a new prescribed random spike train). The dotted vertical lines indicate the limit $T_{\text{max}}$ of almost sure feasibility: for $T\leq T_{\text{max}}$, no unfeasible spike train was observed. For $T>T_{\text{max}}$, the solid line is a logistic regression to the data points.
@@
Note that Fig. 8 suggests that the maximal length $T$ of stably memorizable content grows at least linearly with $K$. (A similar scaling was mathematically proved in a discrete-time setting by Murer and Loeliger (2020).)
6 Small-Jitter Analysis
A necessary condition for temporal stability is that the propagation of spike jitter through the network is damped rather than increased. For small jitter, this can be analyzed by the eigenvalues of a linearized model, as was done in (Banerjee et al., 2008). We are now going to do this as well, and we will find that satisfaction of (8)–(12) (with sufficiently large $\dot{\theta}_{\mathrm{s}}$ and sufficiently small $w_{\mathrm{b}}$ and $\theta_{\mathrm{r}}$) guarantees the suppression of sufficiently small jitter, in agreement with the pertinent remarks in Sec. 4.
6.1 The Math
@@
where we assume (without loss of essential generality) that the influence of $\Delta_{n-n^{\prime}}$ for $n^{\prime}>N$ can be neglected. The derivatives (15) are given by (A.9) in Appendix Appendix A: The Derivatives (15).
@@
6.2 The Use
The eigenvalue condition (20) is necessary, but not sufficient, for temporal stability (with finite threshold noise variance) in the sense of Sec. 3. However, computing the eigenvalues of $\boldsymbol{\Phi}_{N}$ is less time-consuming than, and provides complementary insight to, simulations as in Sections 35.
Complementing the observations in Sec. 4, we experimentally find that satisfying (8)–(12) suffices to satisfy also the eigenvalue condition (20). With the same parameters as for Tables 3 and 4, we evaluate the eigenvalues of $\boldsymbol{\Phi}_{N}$ and report the two largest (in magnitude) $|\varphi_{1}|$ and $|\varphi_{2}|$ in Tables 6 and 6. Every row summarizes 100 repetitions of the same experiment, each with a new random network and a new random prescribed firing score; every row reports the median, the minimum, and the maximum (in magnitude) of these eigenvalues.
Note that the first row of Table 6 demonstrates that dropping the minimal-slope condition (11) makes the network unstable.
Table 5: Effect of the minimum-slope condition (11) on the eigenvalue condition (20), cf. Table 3. Every row summarizes 100 repetitions of the same experiment.